A firm that pays out 65% of its earnings as dividends has an accounting rate of return of 20%.
Its P/E ratio is 10 and its earnings per share is 108 cents.

(i) What is the price per share?

(ii) What is the dividend yield?

(iii) If shares were bought, what would be the payback period? Assume the only return is the dividend.

(iv) What is the net book value per share of the asset investment of the company?

(v) If the risk-adjusted required rate of return is 6%, what would be the NPV per share for buying shares?

(vi) Would you buy shares using AROR or NPV?
Answer #1

(i)P/E Ratio=(Price per share)/(Earning Per Share)

10=(Price per share) /1.08

Price per share=10*1.08=$10.80

(ii) Dividend Yield=(Dividend per share)/(Price per share)

Dividend per share=65%*1.08=$0.70

Dividend Yield=0.70/10.80=0.065=6.5%

(iii) Payback Period in years=Price per share/Dividend per share=10.80/0.70=15.4 years

(iv)Net Book Value per share:

Earning per share =$1.08

Accounting rate of return =20%

(Book value per share )*(Accounting rate of return)=Earning per share

Book Value per share =1.08/20%=$5.40

(v)NPV per share:

Present value of Constant dividend=$0.70/(Required Return)=0.70/0.06=$11.70

Present Price =$10.80

NPV per share=$11.70-$10.80=$0.90

Vi)We should buy the share based on AROR. NPV assumed holding the share in perpetuity . Shares are normally not purchased for holding in perpetuity